# Deblur4DGS: 4D Gaussian Splatting from Blurry Monocular Video

### v2

@@ B Other Regularization Terms
where $\mathbf{M}_{t}$ is the mask obtained by applying SAM2 (Ravi et al. 2024) to the ground truth blurry frame. Besides, we render the 2D tracks $\hat{\mathbf{U}}_{t\rightarrow t^{\prime}}$ from a pair of randomly sampled query time $t$ and target time $t^{\prime}$. We supervise it by the lifted long-range 2D tracks $\mathbf{U}_{t\rightarrow t^{\prime}}$ that are extracted from TAPIR (Doersch et al. 2023), i.e.,
@@ C More Comparison Results
fig. 6 shows some examples of real-world videos captured by the Redmi K50 Ultra smartphone. We evaluate methods on four tasks, i.e., novel-view synthesis, deblurring, frame interpolation, and video stabilization. Per-scene results for novel-view synthesis are summarized in table 7, table 8, table 9 and table 10. Rendering speed comparisons are in table 11. More visual comparisons for novel-view synthesis are in fig. 8 and fig. 9. They show that our Deblur4DGS produces more photo-realistic details and fewer visual artifacts. Besides, table 12, table 14 and table 16 summarize the results for deblurring, frame interpolation, and video stabilization on synthetic videos. table 13, table 15, and table 17 summarize those on real-world videos. Deblur4DGS outperforms state-of-the-art 4D reconstruction methods and has competitive capabilities in comparison with task-specific video processing models trained in a supervised manner. Furthermore, visual results for deblurring in fig. 10 show that compared with 4D reconstruction methods, Deblur4DGS produces sharper contents. We provide some videos in the local HTML file.
@@ D More Ablation Results
Visual results in fig. 5 validate the effectiveness of the proposed regularization terms. Without these regularization terms, notable artifacts appear in dynamic areas. By regularizing the object motion within the exposure time distinguished, $\mathcal{L}_{e}$ improves performance. Besides, $\mathcal{L}_{mfc}$ and $\mathcal{L}_{mrc}$ additionally regularize multi-frame and multi-resolution consistency respectively, helping to alleviate artifacts. In addition, visual results in fig. 7 validate the effectiveness of blur-aware variable (BAV) canonical Gaussians.

### v3

@@ Conclusions
Acknowledgements
This work was supported by the National Natural Science Foundation of China under Grant No. 62371164 and the National Key RD Program of China under Grant No. 2022YFA1004100.
@@ B Other Regularization Terms
where $\mathbf{M}_{t}$ is the mask obtained by applying SAM2 (Ravi et al. 2024) to the ground truth blurry frame. Besides, we render the 2D tracks $\hat{\mathbf{U}}_{t\rightarrow t^{\prime}}$ from a pair of randomly sampled query time $t$ and target time $t^{\prime}$. We supervise it by the lifted long-range 2D tracks $\mathbf{U}_{t\rightarrow t^{\prime}}$ that are extracted from TAPIR (Doersch and et al. 2023), i.e.,
@@
[…]
Figure 8: Visual comparisons with methods that pre-process the blurry video with an image (i.e., Restormer (Zamir et al. 2022)) or video (DSTNet (Pan et al. 2023) and BSSTNet (Zhang and et al. 2024)) deblurring method before performing 4D reconstruction (i.e., Shape-of-Motion (Wang et al. 2024)).
[…]
Figure 9: Example of masks in extreme blurry videos. SAM 2 (Ravi and et al. 2024) may fail to capture small dynamic objects under such extreme blur, as shown in the red box.
@@ C More Comparison Results
fig. 6 shows some examples of real-world videos captured by the Redmi K50 Ultra smartphone. We evaluate methods on four tasks, i.e., novel-view synthesis, deblurring, frame interpolation, and video stabilization. Per-scene results for novel-view synthesis are summarized in table 7, table 8, table 9 and table 10. Rendering speed comparisons are in table 11. More visual comparisons for novel-view synthesis are in fig. 10 and fig. 11. They show that our Deblur4DGS produces more photo-realistic details and fewer visual artifacts. Additionally, the visual comparisons with methods that pre-process the blurry video with an image (i.e., Restormer (Zamir et al. 2022)) or video (DSTNet (Pan et al. 2023) and BSSTNet (Zhang and et al. 2024)) deblurring method before 4D reconstruction (i.e., Shape-of-Motion (Wang et al. 2024)) are in fig. 8, which shows that our Deblur4DGS still performs better.
Besides, table 12, table 14 and table 16 summarize the results for deblurring, frame interpolation, and video stabilization on synthetic videos. table 13, table 15, and table 17 summarize those on real-world videos. Deblur4DGS outperforms state-of-the-art 4D reconstruction methods and has competitive capabilities in comparison with task-specific video processing models trained in a supervised manner. Furthermore, visual results for deblurring in fig. 12 show that compared with 4D reconstruction methods, Deblur4DGS produces sharper contents. Moreover, the visual results of images within an exposure time and the corresponding synthetic blurry image are in fig. 13. It shows that Deblur4DGS successfully synthesize the blur by estimating camera motion and object motion trajectory. We provide some videos at https://deblur4dgs.github.io/.
@@ D More Ablation Results
Visual results in fig. 5 validate the effectiveness of the proposed regularization terms. Without these regularization terms, notable artifacts appear in dynamic areas. By regularizing the object motion within the exposure time distinguished, $\mathcal{L}_{e}$ improves performance. Besides, $\mathcal{L}_{mfc}$ and $\mathcal{L}_{mrc}$ additionally regularize multi-frame and multi-resolution consistency respectively, helping to alleviate artifacts. In addition, visual results in fig. 7 validate the effectiveness of blur-aware variable (BAV) canonical Gaussians. Note that we extract dynamic masks using SAM 2 (Ravi and et al. 2024), which is generally robust enough to produce accurate masks even in blurry videos. However, under extreme blur, it may fail to capture small dynamic objects, as shown in fig. 9.
@@ Methods
[…]
Figure 13: Visual results of images within an exposure time and the synthetic blurry image. The red line is a horizontal reference line. We indicate some regions for easier observation with yellow arrows